# Rethinking Explainable Machine Learning as Applied Statistics

### v4

@@ 7 Related Work
Recently, researchers have begun to take a more statistical approach towards explainable machine learning. For example, Bilodeau et al. (2024) study feature attribution methods via their ability to reject hypotheses about model behavior. Sharma et al. (2024) connect practices in explainable machine learning with p-hacking (Stefan & Schönbrodt, 2023), and Senetaire et al. (2023) formulate feature importance as a statistical inference problem. Cherian & Candès (2024) study fairness auditing from a statistical perspective, and Scholbeck et al. (2024) connect practices in explainable machine learning with sensitivity analysis. To the best of our knowledge, this position paper is the first to make the broad-based case that explainable machine learning is applied statistics for learned functions.
@@ 8 Alternative Views
The most important alternative view to our position is the idea that algorithmic explanations of machine learning systems should be similar to the explanations that humans provide about their own behaviour.777 By attempting to predict machine behavior through postulated intentions, this view takes an “intentional stance” towards machine learning systems. While this approach might be useful to understand human behavior, it is questionable whether it is helpful for algorithms. See, however, Dennett (1987) for a defense of the intentional stance towards early algorithms. According to this perspective, algorithmic explanations would allow humans to understand the behaviour of a machine learning model in the same way that we can understand a human’s behaviour by asking them to explain it (for example, by stating the factors relevant to their decision). While it is rarely explicitly expressed in research papers (see, however, Phillips et al. (2021, Section 8)), this alternative view has had a significant influence on the field of explainable machine learning. In particular, it underlies the perspective that explainable machine learning can attain the same broad-based goals as human explanations, including successful collaboration with humans (Kim, 2015; Lakkaraju et al., 2022), trust (Ribeiro et al., 2016), or general understanding of model behaviour (Hagras, 2018). [^8] We suspect that the choice of the terminology “explanation” in the literature on explainable machine learning contributes to this perspective because it naturally invokes similarities with other forms of explanations, such as human or scientific explanations (Woodward & Ross, 2022).
Related to this is the human-centered perspective on explainable machine learning. Also according to this perspective, explainability is “an essential requirement for people to trust and adopt AI deployed in numerous domains” (Liao & Varshney, 2021). Consequently, explanations should be designed with the various possible users in mind (Liao et al., 2020; Ehsan et al., 2021).
By contrast, we argue that the similarities between human and algorithmic explanations are superficial, that explanations are better understood as statistics, and that future research in explainable machine learning should look for guidance in statistics research rather than elsewhere. We have also argued that explanations are complex technical tools that might mostly be useful to experts. That being said, we agree with many of the propositions in the literature on human-centered explainability. For example, Liao & Varshney (2021) also highlight the importance of intuitive human questions.
@@ 9 Discussion
What about mechanistic interpretability? To be concise, this paper has focused chiefly on post-hoc explanation algorithms. However, we believe that the analogy between explainable machine learning and applied statistics extends to the entire field. As another example, let us briefly outline how it relates to mechanistic interpretability (Saphra & Wiegreffe, 2024).
Mechanistic interpretability aims to reverse-engineer the inner workings of neural networks (Elhage et al., 2021). It analyzes patterns in network activations (Nanda et al., 2023) using a variety of methods, for example probes (Gurnee et al., 2023), or by training sparse auto-encoders (Lieberum et al., 2024; Karvonen et al., 2024). What is the relation between these methods and our work? In a sense, the relationship between applied statistics and mechanistic interpretability is even more apparent than for post-hoc explanations. This is because many of the methods in mechanistic interpretability literally perform statistical research on the dataset of neural network activations. A bit more formally, the activation patterns of a neural network and derived statistics such as probes and sparse auto-encoders fall under the concept of a statistic of a function (Definition 2.3). The interesting difference to post-hoc explanations is that mechanistic interpretability generally aims to find a statistic that can provide insights about a fixed function, whereas a post-hoc method offer a fixed statistic for all functions. Now, the important connection between mechanistic interpretability and our arguments is that once a statistic is chosen, the questions that we should ask are the same as for post-hoc explanations (Section 4): Does it relate to any intuitive concepts? What concrete questions can be answered?999 In some cases, for example, when a probe relates to an interpretable outcome (Karvonen et al., 2024), mechanistic interpretability has found good answers to these questions.
@@ Acknowledgements
We would like to thank Bob Williamson, Julius Adebayo, Suraj Srinivas, Gunnar König and Timo Freiesleben for helpful comments and discussion. This work has been supported by the German Research Foundation through the Cluster of Excellence “Machine Learning - New Perspectives for Science” (EXC 2064/1 number 390727645). Eric Raidl was also funded by the WIN program of the Heidelberg Academy of Sciences and Humanities, financed by the Ministry of Science, Research and the Arts of the State of Baden-Württemberg.

### v5

@@ 8 Related Work
Recently, researchers have begun to take a more statistical approach towards explainable machine learning. For example, Bilodeau et al. (2024) study feature attribution methods via their ability to reject hypotheses about model behavior. Sharma et al. (2024) connect practices in explainable machine learning with p-hacking (Stefan & Schönbrodt, 2023), and Senetaire et al. (2023) formulate feature importance as a statistical inference problem. Cherian & Candès (2024) study fairness auditing from a statistical perspective, and Scholbeck et al. (2024) connect practices in explainable machine learning with sensitivity analysis. To the best of our knowledge, this position paper is the first to make the case that explainable machine learning is applied statistics for learned functions.
An interesting connection exists between our work and the literature on human-centered explainability (Liao et al., 2020; Liao & Varshney, 2021; Ehsan et al., 2021). While we are skeptical that algorithmic explanations can serve as “an essential requirement for people to trust and adopt AI deployed in numerous domains” (Liao & Varshney, 2021), we argue that the literature should focus on the real-world questions that explanations are designed to answer (Section 4). Interestingly, this claim is somewhat similar to the idea that the explanation should focus on the needs of the human users, a tenet in the literature on human-centered explainability (Liao & Varshney, 2021). For example, Figure 1 in Liao et al. (2020) is conceptually similar to the list of intuitive human questions that we provide in Section 4.
@@ 9 Alternative Views
The most important alternative view to our position is the idea that algorithmic explanations of machine learning systems should be similar to the explanations that humans provide about their own behaviour.666 By attempting to predict machine behavior through postulated intentions, this view takes an “intentional stance” towards machine learning systems. While this approach might be useful to understand human behavior, it is questionable whether it is helpful for algorithms. See, however, Dennett (1987) for a defense of the intentional stance towards early algorithms. According to this perspective, algorithmic explanations would allow humans to understand the behaviour of a machine learning model in the same way that we can understand a human’s behaviour by asking them to explain it (for example, by stating the factors relevant to their decision). While it is rarely explicitly expressed in research papers (see, however, Phillips et al. (2021, Section 8)), this alternative view has had a significant influence on the field of explainable machine learning. In particular, it underlies the perspective that explainable machine learning can attain the same broad-based goals as human explanations, including successful collaboration with humans (Kim, 2015; Lakkaraju et al., 2022), trust (Ribeiro et al., 2016), or general understanding of model behaviour (Hagras, 2018). We suspect that the choice of the terminology “explanation” in the literature on explainable machine learning contributes to this perspective because it naturally invokes similarities with other forms of explanations, such as human or scientific explanations (Woodward & Ross, 2022).
By contrast, we argue that the analogy between algorithmic and human explanations is misguided, that explanations are better understood as statistics, and that future research in explainable machine learning should look for inspiration in statistics research rather than elsewhere. We also argue that explanations are complex technical tools that might mostly be useful to experts.
@@ 10 Discussion
Impact Statement
Explainable machine learning is often seen as one of the key components of responsible machine learning. The purpose of our position paper is to encourage discussion about this important topic. As such, we don’t believe that our paper will have negative societal consequences.
@@ Acknowledgements
We would like to thank Julius Adebayo, Timo Freiesleben, Gunnar König, Suraj Srinivas, Bob Williamson and anonymous reviewers for helpful comments and discussion. This work has been supported by the German Research Foundation through the Cluster of Excellence “Machine Learning - New Perspectives for Science” (EXC 2064/1 number 390727645), the CZS Institute for Artificial Intelligence and Law, and the WIN program of the Heidelberg Academy of Sciences and Humanities, financed by the Ministry of Science, Research and the Arts of the State of Baden-Württemberg.